### v2

# Probing Perceptual Constancy in Large Vision-Language Models

Abstract
Perceptual constancy is the ability to maintain stable perceptions of objects despite changes in sensory input, such as variations in distance, angle, or lighting. This ability is crucial for visual understanding in a dynamic world. Here, we explored such ability in current Vision Language Models (VLMs). In this study, we evaluated 155 VLMs using 236 experiments across three domains: color, size, and shape constancy. The experiments included single-image and video adaptations of classic cognitive tasks, along with novel tasks in in-the-wild conditions. We found significant variability in VLM performance across these domains, with model performance in shape constancy clearly dissociated from that of color and size constancy.
1 Introduction
Perceptual constancy is the ability to perceive the properties of environmental objects as stable despite variations in external conditions. While this phenomenon exists across multiple sensory modalities, including auditory and tactile perception, it is particularly well-studied in vision. The human visual system demonstrates remarkable constancy, maintaining stable object perception despite changes in distance, viewing angle, lighting (Epstein, 1977; Walsh and Kulikowski, 1998). This stability allows humans to perceive objects quickly and accurately, even in complex and dynamic scenes, ensuring seamless interaction with the surrounding world.
The study of perceptual constancy dates back to the 19th century, with early theories proposing that it emerges from the brain’s inference process—where stable perception is achieved by integrating empirical knowledge and environmental cues (Von Helmholtz, 1867). Neuroscientific research has since demonstrated that visual information is processed hierarchically, from the retina to the visual cortex (e.g., V1, V4, IT cortex), to maintain perceptual stability across changing conditions (DiCarlo et al., 2012). For example, neurons in the IT cortex consistently encode color regardless of illumination, suggesting that different levels of the visual system progressively compensate for environmental variations to preserve object properties (Tanaka, 1996). This intricate mechanism enables humans to navigate dynamic environments, recognize objects across varying perspectives, and make accurate judgments in complex real-world scenarios. Similarly, instilling AI models with human-like perceptual constancy is crucial for developing robust vision-language models (VLMs) that can perform reliably under diverse and unpredictable conditions. Given the rapid progress of VLMs, perceptual constancy therefore appears to be a key benchmark for assessing their cognitive capabilities and identifying their limitations.
Perceptual constancy is essential for overcoming visual ambiguities in everyday life. Color constancy allows us to recognize objects consistently despite changes in lighting—a critical ability for autonomous systems in varying illumination conditions, such as self-driving cars interpreting traffic signals at different times of day (Land and McCann, 1971). Size constancy ensures that objects appear the same size regardless of distance, facilitating spatial awareness and depth perception. This is particularly important for AI applications in robotics, where accurate size perception is necessary for grasping and object manipulation (Carlson, 2010). Shape constancy enables stable object recognition across different viewpoints, supporting real-world tasks such as facial recognition, medical imaging, and augmented reality applications (Sternberg, 2006). Without these constancy mechanisms, AI systems may struggle with perceptual inconsistencies, leading to unreliable in-the-wild performance that lacks adaptability and precision.
To systematically assess the perceptual constancy capabilities of VLMs, we leveraged the ConstancyBench from the CoreCognition benchmark (Li et al., 2024b), formulating a dataset comprising 253 cognitive experiments. These experiments specifically examine the three primary dimensions of perceptual constancy: color, size, and shape. By testing 155 VLMs on these fundamental perceptual tasks, we aim to uncover the extent to which current models exhibit perceptual stability and where they fall short. Understanding these limitations is essential for advancing AI’s ability to process visual information in a way that aligns more closely with human perception, ultimately enabling more reliable and adaptable AI systems in real-world applications.
Figure 1: Sample Tasks from the Three Evaluation Dimensions of ConstancyBench. Example model performance from GPT-4o is presented.
2 Methods
2.1 Experiment Design
Perceptual constancy refers to the human ability to perceive object properties consistently across varying environmental conditions, here focusing on three key domains: color, size, and shape constancy. These three aspects were selected to capture the fundamental principles of constancy in visual perception, enabling stable object recognition despite changes in lighting, distance, and viewing angle. Below, we provided explanations for each domain.
2.1.1 Color Constancy
Color constancy is an important feature of the human visual system. It allows us to perceive the color of objects consistently under different light conditions. A common example is a white wall appearing in different shades under different lighting conditions. However, the human visual system still perceives it as white rather than the wall’s actual color has changed. This occurs because the visual system can separate an object’s true color from the influence of lighting condition, thereby maintaining stable color perception (Jameson and Hurvich, 1989). Evaluating color constancy can reveal whether VLMs can truly understand an object’s intrinsic color rather than merely relying on color patterns in the training data.
2.1.2 Size Constancy
Size constancy refers to the perception of an object’s size as stable, even when its retinal image changes due to variations in distance (Sperandio and Chouinard, 2015). For example, a distant car appears just as large as a nearby one, despite the difference in retinal projection. This stability is crucial for spatial awareness, depth perception, and navigation. Assessing this phenomenon in VLMs can determine whether they truly grasp the spatial properties of objects in the dynamic environment.
2.1.3 Shape Constancy
Shape constancy allows us to recognize objects as having the same shape, even when viewed from different angles (Rock, 1973). A round plate, for example, may project an elliptical image when seen obliquely, yet we still perceive it as circular. This perceptual stability relies on depth cues, prior experience, and contextual information. Shape constancy is fundamental to object recognition and spatial reasoning, allowing for accurate identification across perspectives. Research suggests humans achieve this by comparing novel views to stored shape representations (Tarr, 1995). Evaluating shape constancy in VLMs can reveal whether they truly understand an object’s three-dimensional form or rely on fixed representations contingent to certain viewpoints.
2.2 Examined Vision Language Models
Recent advances in multi-modal learning have been driven by the unified modeling of visual and textual modalities using transformers (Li et al., 2019; Xu et al., 2023; Tan and Bansal, 2019; Alayrac et al., 2022; Radford et al., 2021). With the rise of large language models (LLMs), state-of-the-art (SOTA) multi-modal LLMs (MLLMs) (Liu et al., 2024; Li et al., 2023b) adopt open-source LLMs (Touvron et al., 2023; Peng et al., 2023; Jiang et al., 2023) and align visual features to the LLM embedding space (Li et al., 2023a; Fu et al., 2023; Wu and Xie, 2024; Xu et al., 2024; Shao et al., 2024; Li et al., 2022; 2025; Brown et al., 2020; Achiam et al., 2023; Bai et al., 2023; Jaech et al., 2024; Zhang et al., 2025; 2024). Progressively, MLLMs have demonstrated competitive performance in complex tasks involving high-level perception and reasoning (Li et al., 2024a; Gemini, 2023; Fu et al., 2023; OpenAI, 2023), such as spatial reasoning (Chen et al., 2024; Cai et al., 2024), character recognition (Mori et al., 1999), scene understanding (Cordts et al., 2016; Wang et al., 2023; Li et al., 2023c; Chen et al., 2017), action recognition (Jhuang et al., 2013; Herath et al., 2017) and prediction (Lan et al., 2014; Kong and Fu, 2022), reaching near-human performance.
We evaluated a total of 155 models for perceptual constancy analysis. Model performances are reported across three domains—color, size, and shape constancy. Each model was tested under a zero-shot setting, generating answers and textual explanations for each experimental prompt. Model size data and architecture type were recorded for correlation analysis.
2.3 Data Sources
The dataset includes six types of data sources: photographs (videos) taken, images (videos) from past classic cognitive psychology experiments, movies (animations), hand-drawn works, and AI-generated images (videos). The distribution is as follows:
  1. 1.
    There are 223 instances of photographs (videos);
  2. 2.
    There are 25 instances of images (videos) from classic cognitive psychology experiments;
  3. 3.
    There are 5 AI-generated images, involving standard 3D geometric shapes only.
Some photographs and video materials have undergone post-processing for experimental purposes.
3 Results
3.1 General Results
Across 155 evaluated Vision-Language Models (VLMs), perceptual constancy performance varies substantially both across and within architectures. Accuracy spans a wide range—from below 0.20 in lightweight open-source models to above 0.90 in frontier systems such as GPT-4o and Gemini 1.5 Pro—highlighting a striking heterogeneity in perceptual stability. This variability underscores that perceptual constancy remains an emergent rather than universal capability: while certain large models exhibit localized robustness under specific transformations, no current architecture demonstrates consistent generalization across all constancy dimensions.
Figure 2: Bar plots show accuracy scores across vision–language models for color ($n=153$), shape ($n=152$), and size ($n=149$) constancy tasks after outlier removal (Color: $0.588\pm 0.185$; Shape: $0.723\pm 0.170$; Size: $0.584\pm 0.123$). Horizontal bars with asterisks indicate statistical significance from post-hoc Tukey HSD tests (* $p<0.001$; ns = not significant, $p>0.05$).
3.2 Domain-wise Differences and Statistical Analysis
Figure 2 summarizes model performance across the three perceptual constancy domains. VLMs demonstrate a clear performance hierarchy, achieving highest accuracy on shape constancy ($M=0.723$, $SD=0.170$), followed by color constancy ($M=0.588$, $SD=0.185$) and size constancy ($M=0.584$, $SD=0.123$). Statistical analysis confirmed these domain-specific differences through a one-way ANOVA revealing a robust main effect of perceptual domain on accuracy ($F(2,451)=36.49$, $p=2.05\times 10^{-15}$, $\eta^{2}=0.139$). Post-hoc Tukey HSD comparisons demonstrated that shape constancy performance significantly exceeded both color ($p<0.001$, Cohen’s $d=0.78$) and size constancy ($p<0.001$, Cohen’s $d=0.89$), while color and size constancy showed statistically equivalent performance ($p=0.976$, Cohen’s $d=0.02$).
These findings suggest shape constancy exploits simple geometric shortcuts, making it the most tractable domain for current VLMs. In contrast, color constancy requires high-dimensional photometric representations, while size constancy demands 3D world representation—explaining why shape dominates while color and size remain computationally challenging across models.
3.3 Relationship Between Model Performance and Model Size
To evaluate how architectural scale influences perceptual constancy, we examined the relationship between model performance and parameter count (Figure 3). Across 151 VLMs for which parameter estimates were available, regression analyses revealed a robust and consistent scaling effect. Overall accuracy increased significantly with model size ($R^{2}=0.2804$, $t=7.62$, $p=2.74\times 10^{-12}$), following the log-linear relation $y=0.1200x+0.4766$ where $x=\log_{10}(\text{param})$. Similar positive correlations were observed in all three perceptual constancy dimensions: color constancy ($R^{2}=0.0973$, $t=4.01$, $p=9.66\times 10^{-5}$, $y=0.0993x+0.4888$), shape constancy ($R^{2}=0.1148$, $t=4.40$, $p=2.08\times 10^{-5}$, $y=0.1017x+0.6171$), and size constancy ($R^{2}=0.3080$, $t=8.14$, $p=1.42\times 10^{-13}$, $y=0.1301x+0.4457$).
These results collectively demonstrate that larger VLMs tend to achieve systematically higher perceptual stability across all constancy domains. The scaling relationships are not only statistically significant but also qualitatively consistent: as model capacity increases, representations become progressively more invariant to color, size, and shape perturbations. Among the three dimensions, the strongest scaling effect was observed in size constancy, indicating that geometric and spatial reasoning benefits most directly from expanded model capacity. In contrast, color constancy showed a weaker but still reliable dependence on size, suggesting that photometric invariance emerges more gradually with scale. Taken together, these findings support a unified scaling law of perceptual stability: the larger the model, the greater its resilience to perceptual variation—a trend that spans all examined dimensions without evidence of domain-specific exceptions.
Figure 3: Relationship between model size and performance. Larger models tend to perform better in perceptual constancy tasks.
3.4 Item Response Theory (IRT) Analysis
To examine the latent structure underlying VLM performance across perceptual constancy tasks, we fitted a two-parameter logistic (2PL) Item Response Theory (IRT) model to all evaluated items. This framework characterizes each item by its discrimination parameter ($a$), which quantifies how sensitively the item differentiates between higher- and lower-ability models, and its difficulty parameter ($b$), which represents the ability level required to achieve a 50% success probability.
Across the three perceptual domains, clear structural distinctions emerged (Figure 4A and Figure 4B). Color items exhibited the highest mean discrimination ($\bar{a}=2.25,\,SD=1.62$) and moderate difficulty ($\bar{b}=-0.58,\,SD=1.32$), indicating that color-related tasks sharply separate stronger from weaker models, but require only moderate ability for success. Shape items showed lower discrimination ($\bar{a}=1.33,\,SD=0.84$) combined with the lowest mean difficulty ($\bar{b}=-1.27,\,SD=1.08$), reflecting tasks that most models could solve reliably and consistently—consistent with their high observed pass rate ($0.76$). In contrast, Size items demonstrated intermediate discrimination ($\bar{a}=1.59,\,SD=1.56$) and the broadest difficulty distribution ($\bar{b}=-0.72,\,SD=1.47$), suggesting greater heterogeneity in spatial reasoning demands and item complexity within this domain.
The global test information curve (Figure 4C) peaks around $\theta\approx 0$, indicating maximal measurement precision for models of average ability, with standard error (SE) below 0.20 across a wide ability range ($1.2<\theta<+1.1$). Together, these results reveal a well-calibrated hierarchy of perceptual challenges: shape constancy tasks anchor the lower end of the difficulty continuum and provide stable measurement of baseline competence, while color and size constancy tasks contribute discriminative power at moderate to high ability levels. This latent structure mirrors behavioral accuracy trends.
Figure 4: Item Response Theory (IRT) analysis of model-level performance. Each point represents a single item parameterized by discrimination ($a$) and difficulty ($b$). Shape-constancy items cluster in regions of lower difficulty and moderate discrimination, whereas color and size items show wider dispersion, reflecting greater variability in perceptual demands and model sensitivity.
4 Discussion
This study introduces ConstancyBench, a large-scale benchmark for evaluating perceptual constancy in 155 Vision-Language Models (VLMs) across three fundamental domains—color, size, and shape constancy. Our results reveal that perceptual constancy is not a uniform emergent ability but a stratified one.
Hierarchical emergence of constancy. The clear dissociation between shape versus color and size constancy supports the existence of a representational hierarchy in how VLMs internalize perceptual stability. Shape constancy—rooted in geometric invariance and low-level topology—appears to emerge from the structural regularities encoded within early convolutional or transformer layers. Such mechanisms correspond to what has been termed minimal constancy: the capacity to preserve stable representations through accumulated perceptual regularities rather than explicit reasoning about context (Bradley, 2008; Buccella, 2021; Buccella and Chemero, 2022). By contrast, color and size constancy require broader contextual integration, including illumination estimation, depth reasoning, and perspective correction—abilities that depend more heavily on multimodal fusion and scene-level abstraction. This hierarchical division parallels the cognitive progression observed in human perception, where geometric stability precedes the development of higher-order contextual invariances.
Scaling dynamics and emergent thresholds. Our scaling analyses reveal that improvements in constancy performance are systematic yet nonlinear. Regression trends indicate that all three domains benefit from increased model capacity, but the relationship is strongest for size constancy, which shows a steep performance gain beyond approximately 10–13 billion parameters. This pattern echoes the “emergent threshold” phenomena widely observed in large language models, where complex behaviors—such as compositional reasoning or in-context adaptation—abruptly appear once sufficient representational depth is achieved. Analogously, perceptual constancy in VLMs may follow a phase-transition dynamic, where representational stability reorganizes once the network’s capacity surpasses a critical complexity threshold. This finding bridges cognitive theories of perceptual maturation with the scaling laws of modern foundation models.
Cognitive interpretation and representational implications. From a cognitive perspective, the observed hierarchy mirrors developmental patterns in human perception: infants acquire shape constancy early through sensorimotor interaction, whereas color and size constancy emerge later as the visual system integrates cross-modal cues. VLMs seem to recapitulate this trajectory at the algorithmic level—progressing from local geometric consistency toward contextual and relational invariance. However, their mechanism of acquisition is fundamentally different: rather than inferential generalization from embodied experience, these models achieve constancy via statistical regularities embedded in massive multimodal datasets. This raises an important conceptual question: do VLMs exhibit functional constancy—genuine invariance to physical transformations—or merely correlational constancy, a byproduct of data coverage and alignment? Addressing this distinction will be key to evaluating whether these systems possess perceptual understanding or only statistical resemblance to it.
Toward mechanistic understanding and future directions. Mechanistically, probing how constancy-related invariances are encoded across visual and cross-attentional layers could reveal whether these patterns reflect bottom-up feature persistence or top-down integration. Architecturally, scaling alone may be insufficient: inductive biases such as explicit 3D priors, lighting normalization, or causal simulation modules may be required to support genuine invariance. Extending ConstancyBench to video, embodied, and generative contexts could further assess how temporal continuity and physical interaction contribute to perceptual stability. Such extensions would enable evaluation of causal perceptual reasoning—the ability not only to maintain perceptual constancy but also to predict and explain variation under changing conditions. Ultimately, perceptual constancy offers a powerful diagnostic lens: it operationalizes the transition from mere perception to world modeling, providing a measurable bridge between cognitive theory and the representational dynamics of modern AI.
5 Conclusion
This work establishes ConstancyBench as the first systematic framework for probing perceptual constancy in large Vision-Language Models (VLMs). Through 236 controlled experiments spanning color, size, and shape constancy, we reveal that current VLMs exhibit a stratified rather than unified form of perceptual stability. Shape constancy emerges robustly even in smaller models, whereas color and size constancy depend strongly on scale and multimodal integration capacity. Our findings uncover a hierarchical pattern of perceptual emergence—one that mirrors developmental trajectories in human perception—and demonstrate that perceptual robustness follows a quantifiable scaling law across modern multimodal systems.
Beyond benchmarking, this study suggests perceptual constancy as a cognitive lens for evaluating the world-modeling fidelity of foundation models. Future work should examine how such invariances arise within model representations, whether through implicit statistical regularities or explicit geometric grounding, and extend these analyses to embodied and dynamic settings. Understanding and mechanistically modeling perceptual constancy will be central to bridging the gap between visual perception and true physical understanding in artificial intelligence.
Appendix A Appendix
A.1 Model accuracy and IRT-derived ability parameters across perceptual dimensions.
Table 1:
Model Accuracy $\theta_{total}$ $\theta_{color}$ $\theta_{shape}$ $\theta_{size}$
o1 0.911 1.707 0.820 0.967 0.929
Qwen2.5-VL-72B-Instruct 0.877 1.394 0.780 0.933 0.897
Eagle-X5-13B-Chat 0.640 1.366 0.950 0.900 0.683
Qwen2.5-VL-72B-Instruct_video 0.877 1.361 0.780 0.900 0.904
qwen-vl-max 0.852 1.356 0.800 0.933 0.881
gemini-1.5-pro 0.788 1.307 0.860 0.800 0.763
gemini-1.5-flash 0.814 1.118 0.820 0.900 0.795
Qwen2-VL-72B-Instruct 0.847 1.083 0.760 0.933 0.859
MiniCPM-Llama3-V-2_5 0.742 1.013 0.840 0.900 0.679
Eagle-X5-34B-Chat 0.619 0.983 0.825 0.800 0.698
gpt-4-turbo 0.847 0.979 0.820 0.900 0.846
gpt-4o 0.843 0.902 0.760 1.000 0.840
LLaVA-NeXT-Video-7B_multi_frame 0.403 0.873 0.857 0.562 0.519
Janus-Pro-7B 0.593 0.858 0.975 0.850 0.604
h2ovl-mississippi-1b 0.716 0.844 0.820 0.900 0.647
xgen-$m$-phi3-dpo-r-v1.5 0.784 0.816 0.800 0.833 0.769
gemini-1.5-flash-8b 0.775 0.801 0.720 0.867 0.776
Llama-3-LongVILA-8B-512Frames 0.500 0.766 0.780 0.533 0.404
Ovis1.5-Gemma2-9B 0.686 0.756 0.940 0.767 0.590
LLaVA-Video-7B-Qwen2_multi_frame 0.750 0.753 0.880 0.933 0.673
LLaVA-NeXT-Video-7B-DPO_multi_frame 0.445 0.732 0.842 0.600 0.500
Kosmos2 0.475 0.728 0.700 0.500 0.400
Mantis-8B-Fuyu 0.555 0.721 0.780 0.714 0.468
InternVL-Chat-V1-2 0.703 0.720 0.740 0.800 0.673
xgen-$m$-phi3-interleave-r-v1.5 0.771 0.712 0.820 0.833 0.744
VideoChat2_stage3_Mistral_7B 0.623 0.690 0.796 0.800 0.538
Llama-3-LongVILA-8B-1024Frames 0.564 0.684 0.880 0.800 0.417
Ovis1.5-Llama3-8B 0.674 0.667 0.780 0.933 0.590
LLaVA-NeXT-Video-32B-Qwen_multi_frame 0.746 0.653 0.720 0.933 0.718
llava_next_mistral_7b 0.665 0.652 0.680 0.867 0.622
LLaVA-Video-7B-Qwen2 0.703 0.610 0.840 0.900 0.622
LLaVA-Video-72B-Qwen2_multi_frame 0.780 0.606 0.740 0.900 0.769
InternVL2_5-78B 0.754 0.581 0.600 0.867 0.782
claude-3-7-sonnet-20250219 0.708 0.571 0.714 0.867 0.757
Llama-3-LongVILA-8B-256Frames 0.542 0.568 0.800 0.733 0.423
InternVL2-8B-MPO-CoT 0.708 0.562 0.740 0.833 0.673
InternVL2-8B-MPO 0.708 0.562 0.740 0.833 0.673
llava-onevision-qwen2-72b-ov-chat-hf 0.708 0.562 0.796 0.733 0.679
llava-onevision-qwen2-72b-ov-hf 0.691 0.561 0.740 0.767 0.673
h2ovl-mississippi-2b 0.513 0.559 0.600 0.633 0.462
Ovis1.6-Gemma2-9B 0.695 0.555 0.740 0.933 0.639
mPLUG-Owl3 0.640 0.543 0.700 0.900 0.571
claude-3-opus-20240229 0.775 0.542 0.860 0.700 0.763
claude-3-5-sonnet-20240620 0.784 0.540 0.760 0.900 0.769
llava-onevision-qwen2-7b-ov-hf 0.686 0.537 0.633 0.733 0.699
deepseek-vl2-small 0.733 0.532 0.720 0.833 0.718
yi-vision-v2 0.699 0.528 0.780 0.800 0.654
Mantis-8B-Idefics2 0.674 0.516 0.740 0.933 0.603
XinYuan-VL-2B-Instruct 0.648 0.511 0.680 0.867 0.596
VideoLLaMA2-72B 0.716 0.489 0.640 0.800 0.724
InternVL-Chat-V1-2-Plus 0.691 0.453 0.660 0.800 0.679
LLaVA-Video-72B-Qwen2 0.733 0.451 0.640 0.867 0.737
video_chat_7b 0.458 0.450 0.587 0.655 0.473
grok-2-vision-1212 0.746 0.427 0.620 0.833 0.769
llava-llama-3-8b 0.606 0.411 0.660 0.800 0.551
Qwen2-VL-7B-Instruct 0.699 0.398 0.700 0.933 0.654
MiniCPM-V-2_6 0.661 0.395 0.840 0.867 0.564
MMAlaya2 0.712 0.377 0.660 0.867 0.699
llava-onevision-qwen2-7b-ov-chat-hf 0.606 0.375 0.633 0.733 0.588
LLaVA-NeXT-Video-32B-Qwen 0.720 0.368 0.640 0.867 0.718
Ovis1.6-Llama3.2-3B 0.648 0.343 0.720 0.800 0.596
Llama-3.2V-11B-cot 0.682 0.336 0.640 0.867 0.660
Aria 0.678 0.329 0.800 0.900 0.596
llava-onevision-qwen2-72b-si-hf 0.665 0.326 0.620 0.800 0.662
Phi-4-multimodal-instruct 0.644 0.300 0.780 0.933 0.545
Qwen2-VL-2B-Instruct 0.661 0.298 0.900 0.833 0.551
llava_next_72b 0.648 0.275 0.680 0.800 0.609
Mantis-8B-siglip-llama3 0.623 0.262 0.660 0.833 0.574
Eagle-X4-13B-Plus 0.551 0.246 0.575 0.950 0.633
llava-onevision-qwen2-7b-si-hf 0.576 0.225 0.653 0.700 0.539
InternVL2_5-26B 0.665 0.220 0.580 0.867 0.654
Llama-3-VILA1.5-8B-Fix 0.665 0.192 0.560 0.767 0.679
pllava-7b 0.551 0.187 0.571 0.750 0.540
hunyuan-vision 0.661 0.183 0.700 0.900 0.603
Phi-3.5-Vision 0.665 0.166 0.660 0.667 0.667
Idefics3-8B-Llama3 0.636 0.153 0.720 0.767 0.583
VILA1.5-40B 0.682 0.152 0.560 0.733 0.712
InternVL-Chat-V1-5 0.665 0.147 0.560 0.867 0.660
VideoLLaMA2-7B 0.470 0.145 0.440 0.567 0.462
VideoChat2_HD_stage4_Mistral_7B_hf 0.475 0.144 0.580 0.667 0.412
Video-LLaVA-7B 0.525 0.125 0.700 0.533 0.468
InternVL2_5-38B 0.606 0.124 0.560 0.700 0.603
llava_next_110b 0.648 0.123 0.600 0.633 0.667
qwen-vl-plus-2025-01-25 0.644 0.121 0.720 0.867 0.596
Qwen2.5-VL-7B-Instruct 0.653 0.116 0.760 0.867 0.577
VideoLLaMA2-7B-16F 0.593 0.109 0.580 0.733 0.571
Qwen2.5-VL-7B-Instruct_video 0.644 0.098 0.740 0.867 0.571
internlm-xcomposer2d5-7b 0.331 0.089 0.641 0.818 0.473
InternVL2-40B 0.682 0.080 0.560 0.933 0.673
InternVL2-8B 0.644 0.080 0.540 0.800 0.647
deepseek-vl-7b-chat 0.525 0.078 0.463 0.900 0.626
video_chat_13b 0.419 0.066 0.457 0.607 0.466
InternVL2-76B 0.669 0.060 0.620 0.933 0.635
InternVL2_5-1B 0.555 0.040 0.540 0.633 0.545
emu2-chat 0.530 0.029 0.480 0.700 0.513
LLaVA-Video-7B-Qwen2-Video-Only_multi_frame 0.504 0.028 0.520 0.700 0.462
Pixtral-12B-2409 0.606 0.027 0.600 0.800 0.571
llava_next_llama3 0.449 0.025 0.525 0.750 0.504
InternVL2-26B 0.648 0.015 0.580 0.767 0.647
VILA1.5-3b 0.513 0.012 0.540 0.700 0.468
SmolVLM 0.602 0.007 0.680 0.667 0.564
LLaVA-NeXT-Video-7B 0.441 0.000 0.804 0.556 0.382
gpt-4o-mini 0.627 -0.035 0.460 0.900 0.628
InternVL2-4B 0.657 -0.043 0.500 0.833 0.673
InternVL2_5-8B 0.551 -0.053 0.520 0.767 0.519
Mini-InternVL-Chat-4B-V1-5 0.602 -0.065 0.520 0.833 0.583
Janus-1.3B 0.445 -0.071 0.450 0.400 0.568
llava_next_interleave_7b_dpo 0.525 -0.096 0.540 0.733 0.481
deepseek-vl2 0.619 -0.102 0.540 0.767 0.615
deepseek-vl2-tiny 0.606 -0.107 0.460 0.800 0.615
internlm-xcomposer2-7b 0.623 -0.109 0.580 0.828 0.614
deepseek-vl-1.3b-chat 0.500 -0.109 0.375 0.600 0.655
Llama-3.2-90B-Vision-Instruct 0.606 -0.115 0.480 0.500 0.667
InternVL2-2B 0.475 -0.124 0.580 0.800 0.378
Llama-3-VILA1.5-8B 0.534 -0.126 0.440 0.633 0.545
InternVL2_5-4B 0.576 -0.133 0.560 0.724 0.558
Mantis-8B-clip-llama3 0.530 -0.142 0.480 0.700 0.513
LLaVA-NeXT-Video-7B-DPO 0.453 -0.142 0.646 0.556 0.418
llava-onevision-qwen2-0.5b-ov-hf 0.517 -0.142 0.440 0.600 0.526
InternVL-Chat-V1-1 0.589 -0.149 0.480 0.733 0.596
LLaVA-Video-7B-Qwen2-Video-Only 0.462 -0.152 0.540 0.633 0.404
claude-3-sonnet-20240229 0.568 -0.161 0.480 0.800 0.555
llava-onevision-qwen2-0.5b-si-hf 0.466 -0.171 0.420 0.500 0.484
Qwen2.5-VL-3B-Instruct 0.542 -0.205 0.380 0.867 0.532
pllava-13b 0.470 -0.205 0.444 0.556 0.524
Llama-3-LongVILA-8B-128Frames 0.542 -0.216 0.540 0.833 0.487
Qwen2.5-VL-3B-Instruct_video 0.542 -0.220 0.400 0.833 0.532
claude-3-haiku-20240307 0.610 -0.223 0.440 0.733 0.641
llava_next_interleave_7b 0.530 -0.238 0.540 0.533 0.526
Chat-UniVi-13B 0.271 -0.269 0.750 0.625 0.522
VideoLLaMA2.1-7B-16F 0.585 -0.296 0.500 0.600 0.609
qwen_chat 0.508 -0.385 0.380 0.700 0.513
VILA1.5-13b 0.534 -0.449 0.460 0.567 0.551
InternVL2_5-2B 0.415 -0.455 0.380 0.667 0.378
Qwen2.5-Omni-7B 0.547 -0.468 0.340 0.733 0.577
Eagle-X5-34B-Plus 0.445 -0.523 0.225 0.500 0.619
Mini-InternVL-Chat-2B-V1-5 0.407 -0.551 0.380 0.800 0.342
Vintern-3B-beta 0.538 -0.587 0.320 0.533 0.609
Eagle-X4-8B-Plus 0.424 -0.601 0.184 0.750 0.561
idefics_9b_instruct 0.453 -0.677 0.383 0.724 0.500
idefics_80b_instruct 0.525 -0.686 0.347 0.633 0.564
JanusFlow-1.3B 0.347 -0.692 0.300 0.450 0.455
VILA1.5-3b-s2 0.500 -0.702 0.320 0.533 0.551
Llama-3.2-11B-Vision-Instruct 0.517 -0.707 0.340 0.567 0.575
OpenFlamingo-9B-vitl-mpt7b 0.483 -0.725 0.340 0.500 0.526
Janus-Pro-1B 0.369 -0.734 0.244 0.700 0.450
InternVL2-1B 0.453 -0.787 0.340 0.633 0.455
Vintern-1B-v2 0.496 -0.787 0.306 0.600 0.538
Eagle-X5-7B 0.403 -0.862 0.200 0.450 0.561
Eagle-X5-13B 0.398 -0.872 0.200 0.500 0.547
Valley2-7b 0.051 -0.882 0.286 0.500 0.267
video_chatgpt-7B 0.225 -0.926 0.271 0.357 0.210
Chat-UniVi 0.318 -0.950 0.340 0.414 0.297
qwen_base 0.140 -1.023 0.400 0.417 0.558
A.2 Item-level discrimination and difficulty parameters estimated from the 2PL IRT model.
Table 2: $a$ = discrimination, $b$ = difficulty, passrate = proportion correct.
Item ID Subtype $a$ $b$ Passrate
400488 Size 8.241 0.204 0.503
400462 Color 6.436 -0.327 0.820
400487 Size 6.178 0.238 0.480
400486 Size 5.615 0.172 0.520
400515 Color 5.434 0.056 0.597
400500 Color 5.430 -0.293 0.793
400502 Color 5.338 -0.260 0.779
400444 Size 5.184 0.261 0.464
400449 Color 4.969 -0.263 0.773
400440 Size 4.904 0.189 0.507
400475 Size 4.855 -0.376 0.818
400458 Size 4.848 0.230 0.487
400441 Size 4.846 0.335 0.423
1732 Size 4.793 -0.055 0.656
400459 Size 4.753 -0.445 0.840
400484 Size 4.676 0.124 0.550
400467 Color 4.646 -0.307 0.788
400507 Size 4.598 0.373 0.388
400504 Color 4.552 0.041 0.603
400442 Size 4.539 0.480 0.336
400482 Size 4.485 0.224 0.490
400513 Size 4.417 0.362 0.405
400453 Color 4.263 -0.055 0.651
400499 Size 4.234 0.098 0.563
400472 Size 4.219 0.064 0.587
400517 Color 4.104 -0.295 0.772
400470 Size 4.023 0.242 0.477
400456 Size 3.995 -0.496 0.838
1740 Shape 3.977 0.056 0.583
400509 Size 3.873 0.224 0.490
1731 Size 3.790 0.290 0.446
400501 Color 3.690 0.440 0.369
1745 Size 3.672 0.259 0.467
400446 Size 3.667 0.253 0.477
400468 Color 3.551 0.351 0.423
1762 Size 3.396 0.108 0.550
400523 Color 3.385 0.395 0.397
400438 Size 3.380 0.468 0.360
1861 Size 3.376 -0.072 0.649
1744 Color 3.213 0.036 0.593
400471 Size 3.212 0.907 0.174
1765 Size 3.142 0.448 0.368
1730 Color 3.074 1.063 0.132
400437 Color 3.034 0.523 0.340
1736 Size 2.875 0.478 0.365
400520 Color 2.857 -0.026 0.616
1776 Size 2.830 1.140 0.121
1826 Shape 2.825 -0.724 0.874
1733 Color 2.806 0.283 0.460
400495 Size 2.754 0.324 0.443
1803 Size 2.662 0.562 0.329
1798 Size 2.641 0.323 0.440
400524 Color 2.622 0.342 0.433
1742 Size 2.616 -0.419 0.772
1812 Size 2.587 0.278 0.461
1852 Size 2.560 0.516 0.351
400477 Size 2.551 0.408 0.404
1882 Color 2.511 -0.629 0.829
1814 Size 2.505 0.687 0.283
400466 Color 2.482 0.591 0.327
1821 Size 2.466 -0.130 0.651
1817 Shape 2.447 -0.177 0.667
1802 Size 2.407 0.334 0.439
400522 Color 2.373 0.497 0.368
1858 Size 2.339 0.776 0.259
1804 Color 2.330 0.452 0.388
400490 Size 2.282 0.225 0.487
1875 Size 2.281 0.260 0.473
1763 Size 2.275 0.405 0.409
1743 Color 2.268 1.018 0.185
1707 Shape 2.177 -0.730 0.834
1800 Size 2.175 0.457 0.385
1250 Shape 2.157 -0.295 0.713
1783 Size 2.104 0.376 0.430
1241 Shape 2.087 -0.339 0.726
1853 Shape 2.040 -0.497 0.763
1756 Color 2.034 0.120 0.533
1784 Size 1.946 -0.146 0.632
1874 Shape 1.908 -1.248 0.914
1210 Size 1.861 0.352 0.444
1747 Size 1.856 0.714 0.309
1248 Size 1.838 0.690 0.321
1813 Shape 1.835 0.252 0.480
1836 Shape 1.786 -0.009 0.576
1223 Size 1.740 0.024 0.574
1240 Size 1.681 -0.143 0.630
1805 Size 1.630 0.711 0.325
400521 Color 1.562 0.993 0.253
400489 Size 1.553 0.901 0.278
400525 Color 1.547 0.999 0.253
1247 Shape 1.512 0.106 0.537
1001208 Size 1.458 -1.811 0.941
400519 Color 1.416 0.807 0.318
1737 Color 1.394 0.893 0.296
1221 Shape 1.378 -0.457 0.699
1901 Color 1.362 -1.141 0.841
1816 Shape 1.357 -0.431 0.682
1208 Size 1.356 -1.368 0.881
1242 Color 1.345 0.272 0.482
1237 Color 1.332 0.495 0.415
1001723 Size 1.310 0.479 0.424
1725 Size 1.216 1.547 0.181
1888 Color 1.200 0.146 0.513
1001724 Color 1.194 -2.524 0.959
1724 Color 1.190 -3.082 0.980
1773 Size 1.174 2.061 0.114
2001723 Size 1.164 0.646 0.388
1894 Size 1.145 0.656 0.382
400436 Size 1.124 1.252 0.250
1883 Size 1.088 1.091 0.289
1791 Size 1.081 0.515 0.423
1746 Color 1.061 0.945 0.322
1212 Shape 1.052 -2.385 0.934
1224 Color 1.039 -1.650 0.867
1001725 Size 1.009 1.826 0.174
1001704 Shape 0.987 -0.860 0.733
1217 Shape 0.967 -1.744 0.861
1222 Color 0.954 -2.588 0.933
1751 Size 0.923 -2.097 0.887
1220 Size 0.915 -1.449 0.815
1779 Size 0.914 -2.191 0.893
1244 Size 0.878 -2.728 0.927
1831 Shape 0.871 0.333 0.473
3001723 Size 0.870 0.126 0.517
1245 Size 0.859 0.956 0.350
400445 Size 0.852 -2.913 0.932
1704 Shape 0.845 -1.695 0.827
1225 Size 0.844 -2.779 0.925
1249 Size 0.841 -2.301 0.890
1723 Size 0.839 0.470 0.449
1230 Size 0.839 -2.388 0.896
1239 Color 0.823 -2.250 0.881
1228 Color 0.820 -1.720 0.825
1209 Size 0.820 -2.427 0.895
1215 Shape 0.814 -2.115 0.867
1001773 Size 0.812 2.859 0.107
1699 Color 0.807 -2.586 0.901
1227 Size 0.789 -1.469 0.787
400478 Size 0.785 -2.164 0.861
1232 Shape 0.779 -3.015 0.926
1698 Color 0.779 -2.635 0.899
1785 Shape 0.766 -2.463 0.882
1702 Size 0.757 -3.188 0.928
1235 Size 0.750 -1.174 0.735
1891 Color 0.742 0.322 0.477
1238 Shape 0.740 -2.622 0.890
1809 Size 0.736 -2.294 0.860
1739 Size 0.732 -2.738 0.894
1728 Size 0.722 -1.973 0.824
1769 Size 0.704 -1.618 0.778
1706 Shape 0.703 -2.218 0.844
1863 Shape 0.702 -1.711 0.789
1734 Size 0.699 -2.748 0.886
400455 Size 0.698 -1.764 0.793
1219 Shape 0.684 -1.674 0.784
1703 Size 0.682 -2.900 0.893
400457 Size 0.682 -3.629 0.934
1788 Size 0.668 -2.468 0.854
1903 Size 0.667 -3.082 0.900
1758 Size 0.667 -3.099 0.901
1793 Size 0.667 -1.854 0.795
1226 Size 0.664 -2.006 0.812
400448 Size 0.661 -2.561 0.860
1246 Color 0.659 -2.559 0.861
1771 Size 0.656 -0.462 0.605
400480 Size 0.651 -2.878 0.882
1700 Color 0.642 -3.123 0.895
1229 Color 0.642 -1.718 0.776
1218 Color 0.640 -2.403 0.843
2001773 Size 0.634 3.240 0.127
1701 Color 0.624 -2.788 0.867
400479 Shape 0.620 -3.110 0.887
1708 Size 0.615 -1.528 0.742
2001725 Size 0.615 1.499 0.313
1887 Size 0.604 -3.012 0.875
1810 Size 0.594 -1.291 0.707
1849 Size 0.593 -2.856 0.860
1774 Size 0.593 -3.036 0.873
1782 Size 0.578 -1.556 0.733
1766 Size 0.563 -1.345 0.705
1873 Size 0.553 -0.680 0.619
1792 Size 0.551 -2.343 0.803
1829 Shape 0.538 -2.936 0.847
1868 Size 0.537 -2.531 0.813
1764 Size 0.531 -0.399 0.578
1906 Size 0.525 -2.247 0.783
1825 Shape 0.522 -2.507 0.807
1881 Color 0.513 -2.324 0.787
400483 Size 0.506 -1.149 0.664
400511 Size 0.506 -2.526 0.800
1711 Size 0.495 -1.912 0.740
1815 Size 0.489 -1.442 0.691
1854 Size 0.483 -3.171 0.840
1234 Color 0.483 -1.318 0.679
1786 Size 0.472 -1.392 0.680
1869 Size 0.469 -2.487 0.781
1847 Size 0.465 -2.165 0.753
1741 Size 0.464 -1.165 0.653
1778 Size 0.455 -2.041 0.737
1845 Size 0.454 -1.819 0.716
1860 Shape 0.444 -1.384 0.671
1781 Size 0.440 -0.895 0.618
1801 Size 0.438 -0.880 0.616
1738 Size 0.437 1.258 0.384
1859 Shape 0.433 -1.459 0.673
1844 Size 0.432 -2.180 0.740
1880 Size 0.425 -1.788 0.702
1885 Size 0.422 -1.469 0.671
1895 Size 0.420 -1.209 0.645
1770 Size 0.417 -0.522 0.576
1797 Size 0.410 -0.126 0.533
1871 Size 0.410 -1.349 0.656
1710 Size 0.409 -1.685 0.687
1897 Size 0.402 1.797 0.342
1808 Size 0.396 -0.699 0.589
1777 Size 0.389 -1.002 0.617
1796 Size 0.387 1.847 0.342
400473 Size 0.385 0.570 0.463
1211 Size 0.379 -2.272 0.726
1787 Size 0.376 0.166 0.503
1780 Size 0.372 -0.227 0.539
1748 Size 0.366 -2.779 0.755
1902 Size 0.365 -0.394 0.555
1856 Size 0.365 -0.529 0.567
1790 Size 0.365 -0.813 0.592
1729 Size 0.363 1.273 0.401
400493 Size 0.362 -2.697 0.747
1806 Size 0.357 0.573 0.466
1840 Size 0.351 0.001 0.517
1705 Size 0.319 -1.680 0.651
400496 Size 0.300 -2.999 0.733
1752 Size 0.300 -3.221 0.747
400460 Size 0.297 -2.894 0.725
400461 Size 0.293 -3.380 0.752
400447 Color 0.248 -2.001 0.642
400498 Size 0.239 -1.672 0.617
A.3 Additional Examples
Figure 5: Additional Sample Tasks: Color Constancy.
Figure 6: Additional Sample Tasks: Color Constancy.
Figure 7: Additional Sample Tasks: Size Constancy.
Figure 8: Additional Sample Tasks: Size Constancy.
Figure 9: Additional Sample Tasks: Shape Constancy.